# Cluster algebras and quantum cohomology rings: A-type

### v1

@@ Abstract
We construct a cluster algebra structure within the quantum cohomology ring of a quiver variety associated with an $A$-type quiver. Specifically, let $Fl:=Fl(N_{1},\ldots,N_{n+1})$ denote a partial flag variety of length $n$, and $QH_{S}^{*}(Fl)[t]:=QH_{S}^{*}(Fl)\otimes\mathbb{C}[t]$ be its equivariant quantum cohomology ring extended by a formal variable $t$, regarded as a $\mathbb{Q}$-algebra. We establish an injective $\mathbb{Q}$-algebra homomorphism from the $A_{n}$-type cluster algebra to the algebra $QH_{S}^{*}(Fl)[t]$. Furthermore, for a general quiver with potential, we propose a framework for constructing a homomorphism from the associated cluster algebra to the quantum cohomology ring of the corresponding quiver variety.
@@ 1.1 Overview
The theory of cluster algebras, developed in [FZ02, FZ03, BFZ05, DWZ08, DWZ10], has revealed deep connections with enumerative geometry through extensive evidence. For instance, Marsh and Rietsch constructed a Landau-Ginzburg B-model of Grassmannian $Gr(r,N)$ where the coordinate ring of the affine Calabi-Yau variety is a cluster algebra [MR20]. Kontsevich and Soibelman established the link between Donaldson-Thomas theory and the cluster transformation in [KS08], while Nagao demonstrated the connection between Donaldson-Thomas theory and cluster algebras [Nag13], proving five fundamental properties of cluster algebras using this relation.
@@
Our work establishes a natural and direct connection between the cluster algebra and the Gromov-Witten theory. We consider a decorated quiver with potential $(\mathbf{Q}=(Q_{f}\subset Q_{0},Q_{1},W),\mathbf{r})$, as defined in Definition 2.1, alongside a stability condition $\theta\in\chi(G)$. Let $Z:=\{dW=0\}$ denote the critical locus of $W$, and let $\mathcal{Z}:=Z\sslash_{\theta}G$ be the corresponding GIT quotient. The quantum cohomology ring of $\mathcal{Z}$, viewed as a $\mathbb{Q}$-algebra, is a center object of our study. Parallelly, a geometric cluster algebra $\mathscr{A}_{\mathbf{Q}}$ can be constructed for the given quiver, see Definitions 2.4-2.6. We propose a construction of an injective $\mathbb{Q}$-algebra homomorphism from the cluster algebra $\mathscr{A}_{\mathbf{Q}}$ to the quantum cohomology ring of $\mathcal{Z}$ in Conjecture 3.8.This homomorphism maps the initial cluster variables to the Chern polynomials of tautological bundles over the variety $\mathcal{Z}$. We explicitly verify this construction for $A$-type quivers, see Theorem 5.1.
@@
This construction is motivated by the study of Seiberg duality conjecture and the insights of [BPZ15]. Seiberg duality explores the relationship between the enumerative theories of two quivers with potentials related by a quiver quiver mutation. Given a decorated quiver with potential $(\mathbf{Q}=(Q_{f}\subset Q_{0},Q_{1},W),\mathbf{r})$, we perform a quiver mutation at a gauge node, yielding another decorated quiver with potential $\mathbf{Q}^{\prime}=((Q_{f}\subset Q_{0},Q_{1}^{\prime},W^{\prime}),\mathbf{r}^{\prime})$. Let $\theta^{\prime}$ denote the character of the gauge group $G^{\prime}$ which is related with $\theta$ via Equation (3.7), and define the variety $\mathcal{Z}^{\prime}:=\{dW^{\prime}=0\}\sslash_{\theta^{\prime}}G^{\prime}$ as the critical locus of the mutated potential function. The Seiberg duality conjecture asserts that Gromov-Witten theories of $\mathcal{Z}$ and $\mathcal{Z}^{\prime}$ are equivalent, see [Hor13, HT07, BPZ15, GLF16] for physics achievements, and [Rua17] for the mathematical formulation. The authors’ previous works, including [Zha21, HZ23] have proved the Seiberg duality conjecture for $A$-types, star-shaped and $D$-type quivers. Similarly, Priddis-Wen-Shoemaker have verified the Seiberg duality conjecture for PXY and PAXY model in [PSW24]. These works focus primarily on genus zero Seiberg duality conjecture and prove transformation formulas of quasimap small $I$-functions of $\mathcal{Z}$ and $\mathcal{Z}^{\prime}$. In this work, we extended the scope to higher-genus Gromov-Witten theory, proving all-genus Seiberg duality conjecture of quivers that are mutation-equivalent to an $A_{n}$-type quiver, see Theorem 4.1.
@@ 1.2 Gromov-Witten theory and Seiberg Duality conjecture
Consider a quiver with potential (QP) $\mathbf{Q}=(Q_{f}\subset Q_{0},Q_{1},W)$, as defined in Definition 2.1. Decorate the quiver with an integer vector $\mathbf{r}=(r_{i})_{i\in Q_{0}}\in\mathbb{Z}_{>0}$ and choose a character $\theta\in\chi(G)$. This setup allows us to define the variety $\mathcal{Z}:=\{dW=0\}\sslash_{\theta}G$ which is the critical locus of the potential function. When $\mathcal{Z}$ is smooth, we can study its Gromov-Witten theory. For the definition of Gromov-Witten theory, see [RT95, RT97, LT98a], with a comprehensive introduction in [CK99]. When the potential is zero, the variety $\mathcal{Z}$ is simply the quiver variety. We refer $\mathcal{Z}$ to the variety associated with the decorated QP $(\mathbf{Q},\mathbf{r})$.
Applying a quiver mutation to the decorated QP as in Definition 2.2 and Equation (2.4), results in a new decorated QP $(\mathbf{Q}^{\prime}=(Q_{f}\subset Q_{0},Q_{1}^{\prime},W^{\prime}),\mathbf{r}^{\prime})$, see Definition 2.2 and Equation (2.4). Choosing a character $\theta^{\prime}$ of the new gauge group $G^{\prime}$ related to the character $\theta$ of $G$ via Equation (3.7), we define the critical locus $\mathcal{Z}^{\prime}$ of $W^{\prime}$. The Seiberg duality conjecture posits that the Gromov-Witten invariants of $\mathcal{Z}$ and $\mathcal{Z}^{\prime}$ are equivalent, see Conjecture 3.3.
In particular, let $QH^{*}_{S}(\mathcal{Z}):=H^{*}_{S}(\mathcal{Z})\otimes\mathbb{Q}[[q]]$ be the equivariant quantum cohomology ring of $\mathcal{Z}$ where $q=(q_{i})_{i\in Q_{0}\backslash Q_{f}}$ are Kähler variables. A genus zero version of the Seiberg duality conjecture asserts that $QH^{*}_{S}(\mathcal{Z})$ and $QH^{*}_{S}(\mathcal{Z}^{\prime})$ are isomorphic, see Conjecture 3.6. Consider the $A_{n}$-type quiver in Figure 5, decorated by integers $N_{1}<\ldots<N_{n+1}$. Let $\Omega_{n}$ denote the set of all QPs mutation-equivalent to the decorated $A_{n}$-type quiver, see Lemma 2.9 and 2.10 for the descriptions of QPs in this set and Lemma 2.15 for the decoration of each QP in $\Omega_{n}$.
@@
Using the localization theorem [Liu11, GP99], we analyze the one-dimensional graphs in $\mathcal{Z}$ and $\mathcal{Z}^{\prime}$ and their contributions to Gromov-Witten invariants. This leads to the proof of the equivalence of all-genus Gromov-Witten invariants of $\mathcal{Z}$ and $\mathcal{Z}^{\prime}$.
@@
$\gamma_{i}\big{|}_{P}=\gamma_{i}^{\prime}\big{|}_{\varphi(P)}\,.$ (1.1)
@@
$\langle\tau_{a_{1}}\gamma_{1},\ldots,\tau_{a_{m}}\gamma_{m}\rangle_{g,m,\beta}^{\mathcal{Z},S}=\langle\tau_{a_{1}}\gamma_{1}^{\prime},\ldots,\tau_{a_{m}}\gamma_{m}^{\prime}\rangle_{g,m,\beta^{\prime}}^{\mathcal{Z}^{\prime},S}.$ (1.2)
@@ 1.3 Cluster algebra conjecture
Given a QP $\mathbf{Q}=(Q_{0},Q_{1},W)$, we can construct a cluster algebra $\mathscr{A}_{\mathbf{Q}}$ as described in Definition 2.2. Denote the initial cluster variables by $\mathbf{x}=(x_{v})_{v\in Q_{0}}$. Then the cluster algebra $\mathscr{A}_{\mathbf{Q}}$ is a sub-algebra in the rational function field $\mathbb{Q}(x_{1},\ldots,x_{n})$, where $n$ is the number of nodes in $Q_{0}$. Let $(\mathbf{Q},\mathbf{x})$ and $(\mathbf{Q}^{\prime},\mathbf{x}^{\prime}=(x_{1}^{\prime},\cdots,x_{n}^{\prime}))$ be two seeds that are related by a seed mutation at node $v\in Q_{0}$, see Definition 2.5. Then the two sets $\mathbf{x}$ and $\mathbf{x}^{\prime}$ coincide except for the variables $x_{v}$ and $x_{v}^{\prime}$, which satisfy the cluster exchange relation:
@@
For the above mentioned QPs $\mathbf{Q}$ and $\mathbf{Q}^{\prime}$, we decorate them by $\mathbf{r}$ and $\mathbf{r}^{\prime}$ such that $r_{u}$ and $r_{u}^{\prime}$ coincide and $r_{v}$ and $r_{v}^{\prime}$ are related by (2.4). Let $\mathcal{Z}$ and $\mathcal{Z}^{\prime}$ be their critical loci which admit a common torus action $S=\prod_{u\in Q_{f}}(\mathbb{C}^{*})^{r_{u}}$ from frozen nodes. We consider the $\mathbb{Q}$-algebra $QH^{*}_{S}(\mathcal{Z})[t]:=QH^{*}_{S}(\mathcal{Z})\otimes\mathbb{Q}[t]$ which we refer to as the quantum cohomology ring of $\mathcal{Z}$ by abuse of terminology. The equivariant Chern polynomials of tautological bundles and quotient bundles $c_{t}^{S}(S_{k})$ and $c_{t}^{S}(Q_{k})$ lie in $QH^{*}_{S}(\mathcal{Z})[t]$ for each node, as detailed in Section 3.2. We conjecture that the quantum product of $c_{t}^{S}(S_{k})$ and $c_{t}^{S}(Q_{k})$ satisfies the quantum cohomological cluster exchange relation, see Conjecture 3.7 which is an analogue of the cluster exchange relation. Therefore, we expect the existence of a $\mathbb{Q}$-algebra homomorphism from $\mathscr{A}_{\mathbf{Q}}$ to $QH_{S}^{*}(\mathcal{Z})[t]$, which sends the initial cluster variables $\mathbf{x}$ to the Chern polynomials of tautological bundles. The challenge is how to define the image of the non-initial cluster variables. We propose an explicit construction in the following Conjecture.
@@
$\psi:\mathscr{A}_{\mathbf{Q}}\rightarrow QH^{*}_{S}(\mathcal{Z})[t]$ (1.3)
@@
  • $\psi(x_{i}^{\prime})=\begin{cases}(-1)^{N_{f}-N_{k}}\xi_{k}^{-1}\prod_{k\rightarrow j}\xi_{j}c_{t}^{S}(Q_{k})\text{ when }N_{f}(k)>N_{a}(k)\\ (-1)^{N_{a}-N_{k}}\xi_{k}^{-1}\prod_{k\rightarrow j}\xi_{j}c_{t}^{S}(Q_{k})\text{ when }N_{f}(k)>N_{a}(k)\end{cases}$ (1.4)
@@
1.4 $A$-type cluster algebra in the quantum cohomology ring of a flag variety
@@
The quantum cohomology ring of a flag variety has been extensively studied and Fontanine has established a ring representation in [CF99]. However, extracting the quantum product of two cohomology classes from the ring representation is still not easy. Our work provides explicit quantum products for a broader range of cohomology classes.
In the proof of Theorem 1.4, we have utilized results of Fontanine-Kim-Sabbah [CFKS08] on the abelian/nonabelian correspondence for Frobenius manifolds, as well as Webb’s work on abelian/nonabelian correspondence for quasimap $I$-functions in [Web23, Web24].
@@ 1.5 Future directions
There are three directions to develop the relation of cluster algebra and enumerative geometry. First, we will prove the cluster algebra conjecture for more quivers. In our upcoming work, we will prove that the cluster algebra Conjecture 1.2 is true for $D$-type quivers. Secondly, we propose that there is also a cluster algebra structure in the quantum K ring, which is on progress. Thirdly, it is natural to consider the quantum difference D-module, and we expect that it admits a quantum cluster algebra structure.

### v2

# Cluster algebra and quantum cohomology ring: the A-type case

@@ Abstract
This work constructs a cluster algebra structure within the quantum cohomology ring of a quiver variety associated with an $A$-type quiver. Specifically, let $Fl:=Fl(N_{1},\ldots,N_{n+1})$ denote a partial flag variety of length $n$, and $QH_{S}^{*}(Fl)[t]:=QH_{S}^{*}(Fl)\otimes\mathbb{C}[t]$ be its equivariant quantum cohomology ring extended by a formal variable $t$, regarded as a $\mathbb{Q}$-algebra. We establish an injective $\mathbb{Q}$-algebra homomorphism from the $A_{n}$-type cluster algebra to the algebra $QH_{S}^{*}(Fl)[t]$. Furthermore, for a general quiver with potential, we propose a framework for constructing a homomorphism from the associated cluster algebra to the quantum cohomology ring of the corresponding quiver variety.
@@ 1.1 Overview
The theory of cluster algebras, developed in [clusteralg:1, clusteralg:2, clusteralg:3, clusterpot1, clusterpot:2], has revealed deep connections with enumerative geometry through extensive evidence. Kontsevich and Soibelman established the link between Donaldson-Thomas theory and the cluster transformation in [kontsevich2008stability], while Nagao demonstrated the connection between Donaldson-Thomas theory and cluster algebras [nagao2013donaldson], proving five fundamental properties of cluster algebras using this relation.
@@
Our work establishes a natural and direct connection between the cluster algebra and the Gromov-Witten theory. We consider a decorated quiver with potential $(\mathbf{Q}=(Q_{f}\subset Q_{0},Q_{1},W),\mathbf{r})$, as defined in Definition 2.1, alongside a stability condition $\theta\in\chi(G)$. Let $Z:=\{dW=0\}$ denote the critical locus of $W$, and let $\mathcal{Z}:=Z\sslash_{\theta}G$ be the corresponding GIT quotient. The quantum cohomology ring of $\mathcal{Z}$, viewed as a $\mathbb{Q}$-algebra, is a central object of our study. Parallelly, a geometric cluster algebra $\mathscr{A}_{\mathbf{Q}}$ can be constructed for the given quiver, see Definitions 2.4-2.6. We propose a construction of an injective $\mathbb{Q}$-algebra homomorphism from the cluster algebra $\mathscr{A}_{\mathbf{Q}}$ to the quantum cohomology ring of $\mathcal{Z}$ in Conjecture 3.8. This homomorphism maps the initial cluster variables to the Chern polynomials of tautological bundles over the variety $\mathcal{Z}$ and noninitial cluster variables to Chern polynomials of various quotient bundles. We explicitly verify this construction for $A$-type quivers, see Theorem 5.1.
@@
This construction is motivated by the study of the Seiberg duality conjecture and the insights of [benini2015cluster]. Seiberg duality explores the relationship between the enumerative theories of two quivers with potentials related by a quiver mutation. Given a decorated quiver with potential $(\mathbf{Q}=(Q_{f}\subset Q_{0},Q_{1},W),\mathbf{r})$, we perform a quiver mutation at a gauge node, yielding another decorated quiver with potential $\mathbf{Q}^{\prime}=((Q_{f}\subset Q_{0},Q_{1}^{\prime},W^{\prime}),\mathbf{r}^{\prime})$. Let $\theta^{\prime}$ denote the character of the gauge group $G^{\prime}$ which is related with $\theta$ via Equation (7), and define the variety $\mathcal{Z}^{\prime}:=\{dW^{\prime}=0\}\sslash_{\theta^{\prime}}G^{\prime}$ as the critical locus of the mutated potential function. The Seiberg duality conjecture asserts that Gromov-Witten theories of $\mathcal{Z}$ and $\mathcal{Z}^{\prime}$ are equivalent, see [Hori, HoriTong, benini2015cluster, gomis2016m2] for physics achievements, and [nonabelianGLSM:YR] for the mathematical formulation. The authors’ previous works, including [zhang2021gromov, hezh2023seiberg] have proved the Seiberg duality conjecture for $A$-types, star-shaped, and $D$-type quivers. Similarly, Priddis-Wen-Shoemaker have verified the Seiberg duality conjecture for PXY and PAXY model in [priddis2024seiberg]. These works focus primarily on genus zero Seiberg duality conjecture and prove transformation formulas of quasimap small $I$-functions of $\mathcal{Z}$ and $\mathcal{Z}^{\prime}$. In this work, we extended the scope to higher-genus Gromov-Witten theory, proving all-genus Seiberg duality conjecture of quivers that are mutation-equivalent to an $A_{n}$-type quiver, see Theorem 4.1.
@@ 1.2 Gromov-Witten theory and Seiberg Duality conjecture
Consider a quiver with potential (QP) $\mathbf{Q}=(Q_{f}\subset Q_{0},Q_{1},W)$, as defined in Definition 2.1. Decorate the quiver with an integer vector $\mathbf{r}=(r_{i})_{i\in Q_{0}}\in\mathbb{Z}_{>0}$ and choose a character $\theta\in\chi(G)$. This setup allows us to define the variety $\mathcal{Z}:=\{dW=0\}\sslash_{\theta}G$ which is the critical locus of the potential function. When $\mathcal{Z}$ is smooth, we can study its Gromov-Witten theory. For the definition of Gromov-Witten theory, see [MR1366548, MR1483992, GW:LT], with a comprehensive introduction in [GW:mirrorsym]. When the potential is zero, the variety $\mathcal{Z}$ is simply the quiver variety. We refer $\mathcal{Z}$ to the variety associated with the decorated QP $(\mathbf{Q},\mathbf{r})$.
Applying a quiver mutation to the decorated QP as in Definition 2.2 and Equation (4), results in a new decorated QP $(\mathbf{Q}^{\prime}=(Q_{f}\subset Q_{0},Q_{1}^{\prime},W^{\prime}),\mathbf{r}^{\prime})$, see Definition 2.2 and Equation (4). Choosing a character $\theta^{\prime}$ of the new gauge group $G^{\prime}$ related to the character $\theta$ of $G$ via Equation (7), we define the critical locus $\mathcal{Z}^{\prime}$ of $W^{\prime}$. The Seiberg duality conjecture posits that the Gromov-Witten invariants of $\mathcal{Z}$ and $\mathcal{Z}^{\prime}$ are equivalent, see Conjecture 3.3. In particular, let $QH^{*}_{S}(\mathcal{Z}):=H^{*}_{S}(\mathcal{Z})\otimes\mathbb{Q}[[q]]$ be the equivariant quantum cohomology ring of $\mathcal{Z}$ where $q=(q_{i})_{i\in Q_{0}\backslash Q_{f}}$ are Kähler variables. A genus zero version of the Seiberg duality conjecture asserts that $QH^{*}_{S}(\mathcal{Z})$ and $QH^{*}_{S}(\mathcal{Z}^{\prime})$ are isomorphic, see Conjecture 3.6.
Consider the $A_{n}$-type quiver in Figure 5, decorated by integers $N_{1}<\ldots<N_{n+1}$. Let $\Omega_{n}$ denote the set of all QPs mutation-equivalent to the decorated $A_{n}$-type quiver, see Lemma 2.9 and 2.10 for the descriptions of QPs in this set and Lemma 2.15 for the decoration of each QP in $\Omega_{n}$.
@@
Using the localization theorem [GW:GPequiv, liu2011localization], we analyze the one-dimensional graphs in $\mathcal{Z}$ and $\mathcal{Z}^{\prime}$ and their contributions to Gromov-Witten invariants. This leads to the proof of the equivalence of all-genus Gromov-Witten invariants of $\mathcal{Z}$ and $\mathcal{Z}^{\prime}$.
@@
$\gamma_{i}\big{|}_{P}=\gamma_{i}^{\prime}\big{|}_{\varphi(P)}\,.$
@@
$\langle\tau_{a_{1}}\gamma_{1},\ldots,\tau_{a_{m}}\gamma_{m}\rangle_{g,m,\beta}^{\mathcal{Z},S}=\langle\tau_{a_{1}}\gamma_{1}^{\prime},\ldots,\tau_{a_{m}}\gamma_{m}^{\prime}\rangle_{g,m,\beta^{\prime}}^{\mathcal{Z}^{\prime},S}.$
@@ 1.3 Cluster algebra conjecture
Let $(\mathbf{Q},\mathbf{x}=(x_{v})_{v\in Q_{0}})$ be a seed which consists of a QP $\mathbf{Q}=(Q_{0},Q_{1},W)$ and a set of algebraically independent variables $\mathbf{x}=(x_{v})_{v\in Q_{0}}$. There is an associated cluster algebra $\mathscr{A}_{\mathbf{Q}}$ as described in Definition 2.2, which is a sub-algebra in the rational function field $\mathbb{Q}(x_{1},\ldots,x_{n})$, where $n$ is the number of nodes in $Q_{0}$. Let $(\mathbf{Q},\mathbf{x})$ and $(\mathbf{Q}^{\prime},\mathbf{x}^{\prime}=(x_{1}^{\prime},\cdots,x_{n}^{\prime}))$ be two seeds that are related by a seed mutation at node $v\in Q_{0}$, then the two sets $\mathbf{x}$ and $\mathbf{x}^{\prime}$ coincide except for the variables $x_{v}$ and $x_{v}^{\prime}$, which satisfy the cluster exchange relation:
@@
For the above mentioned QPs $\mathbf{Q}$ and $\mathbf{Q}^{\prime}$, we decorate them by $\mathbf{r}$ and $\mathbf{r}^{\prime}$ such that $r_{u}$ and $r_{u}^{\prime}$ coincide and $r_{v}$ and $r_{v}^{\prime}$ are related by (4). Let $\mathcal{Z}$ and $\mathcal{Z}^{\prime}$ be their critical loci which admit a common torus action $S=\prod_{u\in Q_{f}}(\mathbb{C}^{*})^{r_{u}}$ from frozen nodes. We consider the $\mathbb{Q}$-algebra $QH^{*}_{S}(\mathcal{Z})[t]:=QH^{*}_{S}(\mathcal{Z})\otimes\mathbb{Q}[t]$ which we refer to as the quantum cohomology ring of $\mathcal{Z}$ by abuse of terminology. The equivariant Chern polynomials of tautological bundles and quotient bundles $c_{t}^{S}(S_{k})$ and $c_{t}^{S}(Q_{k})$ lie in $QH^{*}_{S}(\mathcal{Z})[t]$ for each node, as detailed in Section 3.2. We conjecture that the quantum product of $c_{t}^{S}(S_{k})$ and $c_{t}^{S}(Q_{k})$ satisfies the quantum cohomological cluster exchange relation, see Conjecture 3.7 which is an analogue of the cluster exchange relation. Therefore, we expect the existence of a $\mathbb{Q}$-algebra homomorphism from $\mathscr{A}_{\mathbf{Q}}$ to $QH_{S}^{*}(\mathcal{Z})[t]$, which sends the initial cluster variables $\mathbf{x}$ to the Chern polynomials of tautological bundles. The challenge is how to define the image of the non-initial cluster variables. We propose an explicit construction in the following Conjecture.
@@
$\psi:\mathscr{A}_{\mathbf{Q}}\rightarrow QH^{*}_{S}(\mathcal{Z})[t]$
@@
  2. 2.
    $\psi(x_{i}^{\prime})=\begin{cases}(-1)^{N_{f}-N_{k}}\xi_{k}^{-1}\prod_{k\rightarrow j}\xi_{j}c_{t}^{S}(Q_{k})\text{ when }N_{f}(k)>N_{a}(k)\\ (-1)^{N_{a}-N_{k}}\xi_{k}^{-1}\prod_{k\rightarrow j}\xi_{j}c_{t}^{S}(Q_{k})\text{ when }N_{f}(k)>N_{a}(k)\end{cases}$
  3. 3.
@@
1.4 Cluster algebra conjecture for A-type quiver
@@
Fontanine has established a ring representation in [qcohflag:CF99] for the quantum cohomology ring of a partial flag variety. Our work provides explicit quantum products for a broader range of cohomology classes.
In the proof of Theorem 1.4, we have utilized results of Fontanine-Kim-Sabbah [abelian/nonabelian:CKB] on the abelian/nonabelian correspondence for Frobenius manifolds, as well as Webb’s work on abelian/nonabelian correspondence for quasimap $I$-functions in [abeliannonabelian:Webb, abelianizationlef:Webb].
@@ 1.5 Future directions
There are three directions to develop the relation of cluster algebra and enumerative geometry. First, we will prove the cluster algebra conjecture for more quivers. In our upcoming work, we will prove that the cluster algebra conjecture 1.2 is true for $D$-type quivers. Secondly, we propose that there is also a cluster algebra structure in the quantum K ring, which is in progress. Thirdly, it is natural to consider the quantum difference D-module, and we expect that it admits a quantum cluster algebra structure.